OpenStudy (anonymous):

A solution of acetic acid in water contains 30% by mass of acetic acid (molar mass=60.05g/mol). The density of this solution is 1.0369 g/mL. What is the molarity of acetic acid

OpenStudy (aaronq):

30% acetic acid in 100 mL of solution FInd mass of 100 mL of solution \(\rho=\dfrac{m}{V}\rightarrow m=\rho*V=1.0369 ~g/mL*100 ~mL=103.69 g\) 30 % acetic acid = 0.3*103.69 g=31.107 g \(M=\dfrac{n}{L}=\dfrac{\dfrac{m}{M_r}}{L}=\dfrac{\dfrac{31.107 g}{60.05~g/mol}}{0.1~L}=5.18~M\approx 5.2 ~M\)